# GaussianCube: Structuring Gaussian Splatting using Optimal Transport for 3D Generative Modeling

### v2

@@ 3.1 Representation Construction
Model architecture. Thanks to the spatially structured organization of the proposed GaussianCube, standard 3D convolution is sufficient to effectively extract and aggregate the features of neighboring Gaussians without elaborate designs. We leverage the popular U-Net network for diffusion Nichol and Dhariwal (2021); Dhariwal and Nichol (2021) and simply replace the original 2D convolution layer to their 3D counterparts. The upsampling, downsampling and attention operations are also replaced with corresponding 3D implementations.
Conditioning mechanism. When performing class-conditional diffusion training, we use adaptive group normalization (AdaGN) Dhariwal and Nichol (2021) to inject conditions of class labels $\bm{c}_{\text{cls}}$ into our model, which can be defined as:
$\text{AdaGN}\left(\bm{f}_{i}\right)$ $=\mathrm{GroupNorm}\left(\bm{f}_{i}\right)\cdot(1+\bm{\gamma})+\bm{\beta},$ (4)
where the group-wise scale and shift parameters $\bm{\gamma}$ and $\bm{\beta}$ are estimated to modulate the activations $\{\bm{f}_{i}\}$ in each residual block from the embeddings of both timesteps $t$ and condition $\bm{c}_{\text{cls}}$.
where the condition signal $\bm{c}_{\text{cls}}$ is only needed when training class-conditioned diffusion models. We additionally add supervision on the image level to ensure better rendering quality of generated GaussianCube, which has been shown to effectively enhance the visual quality in previous works Wang et al. (2023); Müller et al. (2023). Specifically, we penalize the discrepancy between the rasterized images $I_{\text{pred}}$ of the predicted GaussianCubes and the ground-truth images $I_{\text{gt}}$:
$\mathcal{L}_{\text{image }}$ $=\mathcal{L}_{\text{pixel}}+\mathcal{L}_{\text{perc }}$ (6)
$=\mathbb{E}_{t,I_{\text{pred }}}\left(\sum_{l}\left\|\Psi^{l}\left(I_{\text{pred}}\right)-\Psi^{l}\left(I_{\text{gt}}\right)\right\|_{2}^{2}\right)$
$+\mathbb{E}_{t,I_{\text{pred}}}\left(\left\|I_{\text{pred}}-I_{\text{gt }}\right\|_{2}^{2}\right),$
where $\Psi^{l}$ is the multi-resolution feature extracted using the pre-trained VGG Simonyan and Zisserman (2014). Benefiting from the efficiency of both rendering speed and memory costs from Gaussian Splatting Kerbl et al. (2023), we are able to render the full image rather than only a small patch as in previous NeRF-based methods Wang et al. (2023); Chen et al. (2023), which facilitates fast training with high-resolution renderings. Our overall training loss can be formulated as:
where $\lambda$ is a balancing weight.
4 Experiments
4.1 Dataset and Metrics
To measure the expressiveness and efficiency of various 3D representations, we fit 100 objects in ShapeNet Car Chang et al. (2015) using each representation and report the PSNR, LPIPS Zhang et al. (2018) and Structural Similarity Index Measure (SSIM) metrics when synthesizing novel views. Furthermore, we conduct experiments of single-category unconditional generation on ShapeNet Chang et al. (2015) Car and Chair. We randomly render 150 views and fit $32\times 32\times 32\times 14$ GaussianCube for each object. To further validate the strong capability of the proposed framework, we also conduct experiments on Omniobject3D Wu et al. (2023b), which is a challenging dataset containing large-vocabulary real-world scanned 3D objects. We fit GaussianCube of the same dimensions as ShapeNet using 100 multi-view renderings for each object. To numerically measure the generation quality, we report the FID Heusel et al. (2017) and KID Bińkowski et al. (2018) scores between 50K renderings of generated samples and 50K ground-truth renderings at the $512\times 512$ resolution.
4.2 Implementation Details
To construct GaussianCube for each object, we perform the proposed densification-constrained fitting for 30K iterations. Since the time complexity of Jonker-Volgenant algorithm Jonker and Volgenant (1988) is $O\left(N_{\text{max}}^{3}\right)$, we opt for an approximate solution to the Optimal Transport problem. This is achieved by dividing the positions of the Gaussians and the voxel grid into four sorted segments and then applying the Jonker-Volgenant solver to each segment individually. We empirically found this approximation successfully strikes a balance between computational efficiency and spatial structure preservation. For the 3D diffusion model, we adopt the ADM U-Net network Nichol and Dhariwal (2021); Dhariwal and Nichol (2021). We perform full attention at the resolution of $8^{3}$ and $4^{3}$ within the network. The timesteps of diffusion models are set to $1,000$ and we train the models using the cosine noise schedule Nichol and Dhariwal (2021) with loss weight $\lambda$ set to $10$. All models are trained on 16 Tesla V100 GPUs with a total batch size of $128$.
Representation Spatially-structured PSNRnormal-$\uparrow$ LPIPSnormal-$\downarrow$ SSIMnormal-$\uparrow$ Rel. Speednormal-$\uparrow$ Params (M)normal-$\downarrow$
@@ 4.2 Implementation Details
Table 2: Quantitative results of representation fitting on ShapeNet Car. $*{}^{*}$ denotes that the implicit feature decoder is shared across different objects.
@@
FID-50Knormal-$\downarrow$ KID-50K(‰)normal-$\downarrow$ FID-50Knormal-$\downarrow$ KID-50K(‰)normal-$\downarrow$ FID-50Knormal-$\downarrow$ KID-50K(‰)normal-$\downarrow$
@@
Table 3: Quantitative results of unconditional generation on ShapeNet Car and Chair Chang et al. (2015) and class-conditioned generation on OmniObject3D Wu et al. (2023b).
@@
Figure 6: Qualitative comparison of class-conditioned 3D generation on large-vocabulary OmniObject3D Wu et al. (2023b). Our model is able to handle diverse distribution with high-fidelity generated samples.
@@ 4.3 Main Results
3D fitting. We first evaluate the representation power of our GaussianCube using 3D object fitting and compare it against previous NeRF-based representations including Triplane and Voxels, which are widely adopted in 3D generative modeling Chan et al. (2022); Wang et al. (2023); Cao et al. (2023); Müller et al. (2023); Tang et al. (2023c). We also include Instant-NGP Müller et al. (2022) and original Gaussian Splatting Kerbl et al. (2023) for reference, although they cannot be directly applied in generative modeling due to their spatially unstructured nature. As shown in Table 2, our GaussianCube outperforms all NeRF-based representations among all metrics. The visualizations in Figure 3 illustrate that GaussianCube can faithfully reconstruct geometry details and intricate textures, which demonstrates its strong capability. Moreover, compared with over $131,000$ Gaussians utilized in the original GS for each object, our GaussianCube only employs 4×4$\times$4 × less Gaussians due to our densification-constrained fitting. This leads to the faster fitting speed of our method and significantly fewer parameters (over 10×10$\times$10 × less than Triplane), which demonstrate its efficiency and compactness. Notably, for 3D generation tasks, NeRF-based methods typically necessitate a shared implicit feature decoder for different objects, which leads to significant decreases sin fitting quality compared to single-object fitting, as shown in Table 2. In contrast, the explicit characteristic of GS allows our GaussianCube to bypass such a shared feature decoder, resulting in no quality gap between single and multiple object fitting.
@@
Single-category unconditional generation. For unconditional generation, we compare our method with the state-of-the-art 3D generation works including 3D-aware GANs Chan et al. (2022); Gao et al. (2022) and Triplane diffusion models Cao et al. (2023). As shown in Table 3, our method surpasses all prior works in terms of both FID and KID scores and sets new records. We also provide visual comparisons in Figure 5, where EG3D and DiffTF tend to generate blurry results with poor geometry, and GET3D fails to provide satisfactory textures. In contrast, our method yields high-fidelity results with authentic geometry and sharp texture details.
@@
Large-vocabulary class-conditioned generation. We also compare class-conditioned generation with DiffTF Cao et al. (2023) on more diverse and challenging OmniObject3D Wu et al. (2023b) dataset. We achieve significantly better FID and KID scores than DiffTF as shown in Table 3, demonstrating the stronger capacity of our method. Visual comparisons in Figure 6 reveal that DiffTF often struggles to create intricate geometry and detailed textures, whereas our method is able to generate objects with complex geometry and realistic textures.
@@
PSNRnormal-$\uparrow$ LPIPSnormal-$\downarrow$ SSIMnormal-$\uparrow$ FID-50Knormal-$\downarrow$ KID-50K(‰)normal-$\downarrow$
@@
Figure 7: Qualitative ablation of representation fitting. Our GaussianCube achieves superior fitting results while maintaining a spatial structure.
@@
Figure 8: Ablation study of the Gaussian structuralization methods and 3D generation. For visualization of Gaussian structuralization in (a), we map the coordinates of the corresponding voxel grid of each Gaussians to RGB values to visualize the organization. Our Optimal Transport solution yields smooth transit among Gaussians, indicating a coherent global structure, whereas Nearest Neighbor Transport leads to obvious discontinuities. Our OT-based solution also results in the best generation quality shown in (b).
@@ 4.4 Ablation Study
We first examine the key factors in representation construction. To spatially structure the Gaussians, a simplest approach would be anchoring the positions of Gaussians to a predefined voxel grid while omitting densification and pruning, which leads to severe failure when fitting the objects as shown in Figure 7. Even by introducing learnable offsets to the voxel grid, the complex geometry and detailed textures still can not be well captured, as shown in Figure 7. We observe that the offsets are typically too small to effectively lead the Gaussians close to the object surfaces, which demonstrates the necessity of performing densification and pruning during object fitting. Based on these insights, we do not organize the Gaussians during the fitting stage. Instead, we only maintain a constant number of Gaussians using densification-constrained fitting and post-process the Gaussians into a spatially structured representation. Our GaussianCube can capture both complex geometry and intricate details as illustrated in Figure 7. The numerical comparison in Table 4 also demonstrates the superior fitting quality of our GaussianCube.
We also evaluate how the representation affects 3D generative modeling in Table 4 and Figure 8. Limited by the poor fitting quality, performing diffusion modeling on voxel grid with learnable offsets leads to blurry generation results as shown in Figure 8. To validate the importance of organizing Gaussians via OT, we compare with the organization based on nearest neighbor transport between positions of Gaussians and centers of voxel grid. We linearly map each Gaussian’s corresponding coordinates of voxel grid to RGB color to visualize the different organizations. As shown in Figure 8 (a), our proposed OT approach results in smooth color transitions, indicating that our method successfully preserves the spatial correspondence. However, nearest neighbor transport does not consider global structure, which leads to abrupt color transitions. Notably, since our OT-based organization considers the global spatial coherence, both the quantitative results in Table 4 and visual comparisons Figure 8 indicate that our structured arrangement facilitates generative modeling by alleviating its complexity, successfully leading to superior generation quality.
@@ 5 Conclusion
We have presented GaussianCube, a novel representation crafted for 3D generative models. We address the unstructured nature of Gaussian Splatting and unleash its potential for 3D generative modeling. Firstly, we fit each 3D object with a constant number of Gaussians by our proposed densification-constrained fitting algorithm. Furthermore, we organize the obtained Gaussians into a spatially structured representation by solving the Optimal Transport problem between the positions of Gaussians and the predefined voxel grid. The proposed GaussianCube is expressive, efficient and with spatially coherent structure, providing a strong 3D representation alternative for 3D generation. We train 3D diffusion models to perform generative modeling using GaussianCube, and achieve state-of-the-art generation quality on the evaluated datasets without elaborate network or training algorithm design. This demonstrates the promise of GaussianCube to be a versatile and powerful 3D representation for 3D generation.

### v3

@@ 3.1 Representation Construction
Model architecture. Thanks to the spatially structured organization of the proposed GaussianCube, standard 3D convolution is sufficient to effectively extract and aggregate the features of neighboring Gaussians without elaborate designs. We leverage the standard U-Net network for diffusion [38, 15] and simply replace the original 2D operators including convolution, attention, upsampling and downsampling with their 3D counterparts.
Conditioning mechanism. Our model supports a variety of condition signals to control the generation process. When performing class-conditioned diffusion modeling, we employ adaptive group normalization (AdaGN) [15] to inject the class labels into our model. For image-conditioned digital avatar creation, we leverage a pretrained vision transformer [6] to encode the conditional image into a sequence of feature tokens. We subsequently adopt cross-attention to make the model learn the correspondence between 3D activations and 2D image feature tokens following [5]. We also leverage cross-attention as our condition mechanism when creating 3D objects from text, similar to previous text-to-image diffusion models [44].
[…]
@@
Table 2: Comparison with prior 3D representations of spatial structure, fitting quality, relative fitting speed (Rel. Speed) and parameter sizes on ShapeNet Car. denotes that the implicit feature decoder is shared across different objects. All methods are evaluated at 30K iterations.
@@
where the condition signal $\bm{c}_{\text{cls}}$ is only needed when training conditional diffusion models. We additionally impose image-level supervision to improve the rendering quality of generated GaussianCube, which has been demonstrated to effectively enhance the perceptual details in previous works [55, 36]. Specifically, we penalize the discrepancy between the rasterized images $I_{\text{pred}}$ of the predicted GaussianCubes and the ground-truth images $I_{\text{gt}}$ using:
$\mathcal{L}_{\text{image }}=\mathbb{E}_{t,I_{\text{pred }}}\left(\sum_{l}\left\|\Psi^{l}\left(I_{\text{pred}}\right)-\Psi^{l}\left(I_{\text{gt}}\right)\right\|_{2}^{2}\right)+\mathbb{E}_{t,I_{\text{pred}}}\left(\left\|I_{\text{pred}}-I_{\text{gt }}\right\|_{2}\right),$ (4)
where $\Psi^{l}$ is the multi-resolution feature extracted using the pre-trained VGG [47]. Benefiting from the efficiency of both rendering speed and memory costs from GS [28], we are able to perform fast training with high-resolution renderings. Our overall training loss can be formulated as:
where $\lambda$ is a balancing weight.
@@
[…]
@@
Table 3: Quantitative results of unconditional generation on ShapeNet Car and Chair [9] and class-conditioned generation on OmniObject3D [59].
@@
Figure 6: Qualitative comparison of class-conditioned 3D generation on large-vocabulary OmniObject3D [59]. Our model is able to handle diverse distribution with semantically accurate results.
4 Experiments
4.1 Dataset and Metrics
To measure the expressiveness and efficiency of various 3D representations, we fit 100 objects in ShapeNet Car [9] using each representation and report the PSNR, LPIPS [68] and Structural Similarity Index Measure (SSIM) metrics when synthesizing novel views. Furthermore, we conduct experiments of single-category unconditional generation on ShapeNet [9] Car and Chair, and class-conditioned generation on real-world scanned dataset OmniObject3D [59]. We compute the FID [22] and KID [2] scores between 50K generated renderings and 50K ground-truth renderings. For image-conditioned digital avatar generation, we utilize the synthetic avatar dataset [57], which comprises highly-detailed 3D avatars created by synthetic pipeline. We assess the generation quality of 5K rendering from 500 test avatars and additionally include cosine similarity of identity embedding [13] (CSIM) to measure the ID preservation. The experiments of text-to-3D generation are performed on the large-scale challenging Objaverse dataset [12]. We numerically evaluate the text alignment quality using CLIP score [43] of 300 test prompts. All images are rendered with $512\times 512$ resolution. For more details of data, please refer to Section A.1.
[…]
Rodin w/o 2D SR 18.80 0.2842 0.7439 0.6594 32.07 24.78
Rodin 18.59 0.2821 0.7373 0.6466 20.02 9.24
Ours 21.87 0.1768 0.7703 0.7821 8.32 2.67
Table 4: Quantitative results of digital avatar creation conditioned on single portrait image.
[…]
Figure 7: Qualitative comparison of 3D avatar creation conditioned on single frontal portraits.
4.2 Implementation Details
To construct GaussianCube for each object, we perform the proposed densification-constrained fitting for 30K iterations, where $N_{\text{max}}$ is set to 32,768. After OT-based structuralization, we obtain $32\times 32\times 32\times 14$ GaussianCube for each object. For the 3D diffusion model, we adopt the ADM U-Net network [38, 15]. We perform full attention at the resolution of $8^{3}$ and $4^{3}$ within the network. The timesteps of diffusion models are set to $1,000$ and we train the models using the cosine noise schedule [38] with loss weight $\lambda$ set to $10$. For more training details, please refer to Section A.1.
@@ 4.3 Main Results
3D fitting. We first evaluate our representation capability of object fitting against previous NeRF-based representations including Voxels [53] with similar parameter sizes and Triplane [8], which are widely adopted in previous 3D generation works [8, 55, 5, 36, 53]. We also include Instant-NGP [37] and original Gaussian Splatting [28] for reference despite their unsuitability for generative modeling due to their unstructured spatial nature. As shown in Table 2, our GaussianCube outperforms all NeRF-based representations among all metrics. Figure 3 illustrates that GaussianCube can faithfully reconstruct geometry details and intricate textures. Moreover, we achieve such high-quality fitting with minimal parameters due to the densification-constrained fitting, showcasing our compactness. Notably, the shared implicit feature decoder in the multi-object fitting of NeRF-based methods leads to significant decreases in quality compared to single-object fitting as evidenced in Table 2. While the fully explicit nature of GS results in no quality gap between single and multiple object fitting.
DreamGaussian VolumeDiffusion Shap-E LGM Ours
[…]
Inference Time (s)$\downarrow$ $\sim 120$ 5 7 2 5
Table 5: Quantitative results of text-to-3D creation. Inference time is measured on a single A100 GPU. While Shape-E, LGM achieve comparable CLIP scores as ours, they either utilize millions of training data or leverage 2D diffusion prior.
[…]
Figure 8: Qualitative comparison of text-to-3D generation on Objaverse [12]. Our model is able to generate high-quality samples according to the given text prompts.
@@
Single-category unconditional generation. For unconditional generation, we compare our method with the state-of-the-art 3D generation works including 3D-aware GANs [8, 17] and Triplane diffusion models [5]. As shown in Table 3, our method surpasses all prior works in terms of both FID and KID scores and sets new records. We provide visual comparisons in Figure 5, where EG3D and DiffTF tend to generate blurry results with poor geometry, and GET3D fails to provide satisfactory textures. In contrast, our method yields high-fidelity results with authentic geometry and sharp textures.
@@
Large-vocabulary class-conditioned generation. We also compare class-conditioned generation with DiffTF [5] on more diverse and challenging OmniObject3D [59] dataset. We achieve significantly better FID and KID scores than DiffTF as shown in Table 3. Visual comparisons in Figure 6 reveal that DiffTF often struggles to create intricate geometry and detailed textures, whereas our method is able to generate objects with complex geometry and realistic textures.
Image-conditioned avatar generation. For 3D avatar generation conditioned on a single reference image, we compare our method with state-of-the-art Triplane diffusion models, Rodin [44]. Our model surpasses Rodin among all evaluated metrics as shown in Table 4. Although Rodin utilizes a 2D refiner [56] to boost the visual quality of facial areas, which significantly compromises 3D consistency. Our model still outperforms it by direct real-3D generation. Results in Figure 7 demonstrate that our model faithfully preserves the identity, expression and accessories of the references with rich details, while Rodin struggles to provide satisfactory results even using 2D refinement.
@@
[…]
@@
Figure 9: Qualitative ablation of representation fitting.
@@
Figure 10: Visual ablation of the Gaussian organization methods and 3D generation. For visualization of Gaussian structuralization in (a), we map the coordinates of the corresponding voxel of each Gaussians to RGB values to visualize the organization. Our OT-based solution also results in the best generation quality shown in (b).
Text-to-3D generation. We compare text-to-3D generation with prior arts including diffusion models [26, 53], optimization-based method [50] and feed-forward method [51]. Our model achieves competitive text-3D alignment results as shown in Table 5. The visual comparison in Figure 8 shows that our model is able to create high-quality samples aligning with text prompts in 5 seconds. DreamGaussian tends to create over-saturated results and suffers from Janus problem. VolumeDiffusion produces unsatisfactory textures with poor text alignment. Shap-E can produce semantically accurate results but struggles to generate complex geometry. LGM reconstructs 3D Gaussians from multi-view images generated by text-conditioned multi-view diffusion pipeline [45], whereas the inconsistency [51] of the generated multi-views often results in inaccurate geometric reconstruction.
@@ 4.4 Ablation Study
We first examine the key factors in representation construction on ShapeNet Car. To spatially structure the Gaussians, a simplistic approach would be anchoring the positions of Gaussians to a predefined voxel grid while omitting densification and pruning, which leads to severe failure when fitting the objects as shown in Figure 9. Even by introducing learnable offsets to the voxel grid, the results still lack details. We observe the offsets are typically too small to effectively lead the Gaussians close to the object surfaces, which indicates the importance of densification in the fitting process. Instead, GaussianCube can capture both complex geometry and intricate details as shown in Figure 9. The numerical comparison in Table 6 also demonstrates the superior fitting quality of GaussianCube.
We also evaluate how the representation affects 3D generative modeling on ShapeNet Car as shown in Table 6 and Figure 10. Limited by the poor fitting quality, performing diffusion modeling on voxel grid with learnable offsets leads to blurry generation results as shown in Figure 10. To validate the importance of organizing Gaussians via Optimal Transport (OT), we compare with the organization based on nearest neighbor transport. We linearly map each Gaussian’s corresponding coordinates of voxel to RGB color to visualize different organizations. As shown in Figure 10 (a), our proposed OT approach yields smooth color transitions, indicating that our method successfully preserves the spatial correspondence. However, nearest neighbor results in abrupt color transitions due to their disregard for global structure. Both the quantitative results in Table 6 and visual comparisons Figure 10 indicate that our globally structured arrangement facilitates generative modeling by alleviating its complexity, successfully leading to superior generation quality.
@@ 5 Conclusion
We have presented GaussianCube, a structured and explicit radiance representation crafted for 3D generative models. We begin by fitting each 3D object with a constant number of Gaussians using our proposed densification-constrained fitting algorithm. We further organize the obtained Gaussians into a spatially structured representation by solving the Optimal Transport between the positions of Gaussians and the predefined voxel grid. The proposed GaussianCube is spatially structured, allowing to use standard 3D U-Net for diffusion modeling without elaborate designs. Moreover, GaussianCube can achieve high-quality fitting using much fewer parameters compared to prior works of similar quality, which further eases the difficulty of generative modeling. Our 3D diffusion models equipped with GaussianCube achieve state-of-the-art generation quality on the evaluated datasets, underscoring its potential of GaussianCube as a versatile and powerful radiance representation for 3D generation.
Appendix A Appendix
A.1 Additional Implementation Details
Dataset preparation. We conduct experiments on ShapeNet Car [9], ShapeNet Chair [9], OmniObject3D [59], Synthetic Avatar [57] and Objaverse [12] datasets. For each dataset, we report the total number of objects used for training, the number of views rendered per object for GaussianCube fitting and the distribution of camera poses used for rendering in Table 7. For the Objaverse dataset, we excluded low-quality objects, such as those without textures or with defective reconstructions following [53]. We also report the object bounding box $\bm{b}$ in the world coordinate system of each dataset in Table 7, which is used to construct the predefined voxel grid within $[-\bm{b},\bm{b}]^{3}$ during OT-based Gaussian structuralization.
Representation construction. We set $N_{\text{max}}$ to 32768 and $C$ to 14 omitting the view-dependent spherical harmonics. This simplification appears to have a negligible impact on object fitting while concurrently reducing the data dimension, thereby alleviating the difficulty of diffusion modeling. During our densification-constrained fitting procedure, we primarily follow the hyper-parameters in original Gaussian Splatting [28]. For OT-based Gaussian structuralization, we adopt an approximate solution for the OT problem due to the $O\left(N_{\text{max}}^{3}\right)$ time complexity of Jonker-Volgenant algorithm [25]. This is achieved by dividing the positions of the Gaussians and the voxel grid into four sorted segments and then applying the Jonker-Volgenant solver to each segment individually. We empirically find this approximation successfully strikes a balance between computational efficiency and spatial structure preservation. The proposed densification-constrained fitting takes around $2.67$ minutes for each object of 30K iterations and the OT-based voxelization takes around $2$ minutes which can be run on CPU in parallel.
3D Diffusion. To train the 3D diffusion model, we initially compute the instance-wise statistics of mean $\bar{\bm{\mu}}\in\mathbb{R}^{N_{v}\times N_{v}\times N_{v}\times C}$ and standard deviation $\bar{\bm{\sigma}}\in\mathbb{R}^{N_{v}\times N_{v}\times N_{v}\times C}$, from the GaussianCubes of each training dataset respectively. These statistical measures are then utilized to normalize the training data. For our 3D diffusion model architecture, we adopt the ADM-UNet from [15] and replace the convolution, upsampling, downsampling and attention operations with 3D implementations. We train our model using AdamW optimizer [31], and apply exponential moving average (EMA) with a rate of 0.9999 during training. For unconditional generation on ShapeNet, we train the model with a base learning rate $55e-$ 5 for 850K iterations and then decay the learning rate to $65e-$ 6 for another 150K iterations. For 3D digital avatar creation from a single portrait image, we adopt the pretrained DINO ViT-B/16 [6] to encode the $512\times 512$ conditional images into $1025\times 768$ conditional feature tokens. For text-to-3D creation, we take CLIP-L/14 [43] to encode the text prompts into $77\times 768$ conditional feature tokens. We provide more detailed configurations of the model architectures, diffusion training and inference for each dataset in Table 8.
Implementation of Gaussian organization visualization in  Figure 10 (a). For the $i$-th Guassian, we obtain its corresponding voxel grid centers $\bm{x}_{k}\in\mathbb{R}^{3}$ according to Optimal Transport plan $\mathbf{T}^{*}$ (i.e., $\mathbf{T}^{*}_{ik}=1$) as illustrated in Section 3.1. To visualize the coordinates of $\bm{x}_{k}$, we map them to RGB color $\bm{C}_{k}\in\mathbb{R}^{3}$ using:
$\bm{C}_{k}=\frac{(\bm{x}_{k}+\bm{b})}{2\bm{b}}\times\bm{255},$ (6)
where $\bm{b}$ is the bounding box in the world coordinate system. The resultant point cloud like visualizations are shown in Figure 10 (a), where smooth color transitions indicate coherent spatial correspondence preservation.
A.2 Additional Ablation Study and Analysis
Ablation of $N_{\text{max}}$ in densification-constrained fitting. We conduct experiments to evaluate how $N_{\text{max}}$ affects fitting on ShapeNet Car. The results in Table 9 indicate that there is a clear trend where increasing $N_{\text{max}}$ leads to improved fitting accuracy. However, a larger $N_{\text{max}}$ also incurs higher computational costs during diffusion training. Therefore, we set $N_{\text{max}}$ to 32,768 to strike a balance between high-quality fitting and computational efficiency.
Ablation of classifier-free guidance in class-conditioned generation. We study how classifier-free guidance (CFG) impacts our generation quality when inference class-conditioned diffusion models. We report the FID and KID metrics in Table 10 under different CFG scales.
Visualization of intermediate results in the denoising process. During inference, our model starts from Gaussian noise and progressively denoises to yield the high-quality GaussianCube. We present visualizations of the intermediate renderings $\bm{y}_{t}$ at various timesteps $t\in[0,T]$ throughout the denoising process, offering a detailed insight into the GaussianCube diffusion procedure. As illustrated in Figure 11, our model first establishes the global structure and then incrementally enhances the details, which is similar to previous 3D diffusion models [55, 46].
Dataset # Objects # Views per object Rotation Angle Elevation Angle Bounding Box
ShapeNet Car 7,462 150 $[0,2\pi]$ $[\frac{1}{6}\pi,\frac{1}{2}\pi]$ 0.45
ShapeNet Chair 6,775 150 $[0,2\pi]$ $[\frac{1}{6}\pi,\frac{1}{2}\pi]$ 0.35
OmniObject3D 5,795 100 $[0,2\pi]$ $[0,\frac{1}{2}\pi]$ 1.0
Sythetic Avatar 98,000 300 $[0,2\pi]$ $[\frac{1}{6}\pi,\frac{2}{3}\pi]$ 40.0
Objaverse 125,653 150 $[0,2\pi]$ $[0,\frac{2}{3}\pi]$ 0.5
Table 7: Details of each dataset.
ShapeNet Car ShapeNet Car OmniObject3D Synthetic Avatar Objaverse
Diffusion steps 1,000 1,000 1,000 1,000 1,000
Noise schedule Cosine Cosine Cosine Cosine Cosine
NFEs 300 300 300 250 44
Inferece sampler DPM-solver [32] DPM-solver [32] DPM-solver [32] DPM-solver [32] DPM-solver [32]
CFG scale - - 2.0 1.3 3.5
Model size 82M 82M 82M 339M 339M
Channels 64 64 64 128 128
Channel mult (1,2,3,4) (1,2,3,4) (1,2,3,4) (1,2,3,4) (1,2,3,4)
Num res blocks 3 3 3 3 3
Attn resolutions (8, 4) (8, 4) (8, 4) (8, 4) (8, 4)
Num head channels 64 64 64 64 64
Dropout 0 0 0 0 0
Scale shift norm True True True True True
Training steps 1, [^0] K 1, [^0] K 700K 1,200K 1,800K
Training GPUs 16 16 16 16 32
Batch size 128 128 128 128 256
Base lr $55e-$ 5 $55e-$ 5 $55e-$ 5 $55e-$ 5 $55e-$ 5
Lr decay steps 850K 850K - - -
Table 8: Detailed configuration of model architecture, diffusion training and inference on each dataset.
[…]
4096 16 32.56 0.0547 0.9765
13824 24 34.32 0.0396 0.9842
32768 32 34.94 0.0347 0.9863
110592 48 35.29 0.0307 0.9874
262144 64 35.34 0.0301 0.9875
Table 9: Quantitative ablation of $N_{\text{max}}$ in densification-constrained fitting. We set $N_{\text{max}}$ to 32,768 in this paper.
Scale w/o CFG 1.3 1.5 2.0 3.0 6.0
[…]
KID-50K(‰$\downarrow$ 4.01 3.12 3.00 2.78 3.17 14.36
Table 10: Quantitative ablation of CFG scale in the class-conditioned generation of OmniObject3D [59].
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Visualization of generation results in intermediate diffusion timesteps.
Figure 12: Visualization of nearest neighbor search on ShapeNet Car and Chair.
[…]
Figure 13: Failure cases.
Nearest neighbors analysis. We perform nearest neighbor search of some unconditionally generated samples in the paper according to the similarity of pretrained CLIP [43] features. The results in Figure 12 demonstrate that our model is capable of generating novel geometry and textures rather than simply memorizing the training data.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Reference Rodin [55] Ours
Figure 14: Additional qualitative comparison of 3D avatars creation conditioned on single in-the-wild portraits.
[…]
Figure 15: Qualitative comparison generated digital avatars conditioned on synthetic portraits.
[…]
Figure 16: Additional qualitative comparison of text-to-3D generation on Objaverse [12]. Our model is capable of creating high-quality samples following input text prompts.
Figure 17: Additional results of text-to-3D generation.
[…]
Figure 18: Variation of text-to-3D generation. Our model is able to generate diverse results conditioned on the same text prompt.
[…]
Figure 19: Example of text-guided 3D editing.
Figure 20: Additional generated samples on ShapeNet Car.
Figure 21: Additional generated samples on ShapeNet Chair.
Figure 22: Additional generated samples on OmniObject3D.
A.3 Additional Visual Results
For 3D avatar generation, while trained on synthetic dataset, our model is capable of generalizing to in-the-wild portrait input. We provide more visual comparison of 3D avatar creation conditioned on in-the-wild portraits with Rodin [55] in Figure 14. We also include additional comparison conditioned on synthetic input from our test in Figure 15. Our model can faithfully retain the identity of the reference portrait and is able to provide high-fidelity results with rich details, e.g. hair, glasses and clothing. Although utilizing a pretrained 2D super-resolution module which significantly compromises 3D consistency, Rodin struggles to follow the conditional images and fails to produce detailed textures in non-facial areas e.g. clothing and hair.
We include additional qualitative comparison and generated samples of text-to-3D generation in Figure 16 and Figure 17 respectively. Our model yields samples with better visual quality, and is capable of handling challenging prompts. The results in Figure 18 show the generation diversity of our results given the same text prompt. Our model is also capable of performing text-guided editing of generated objects by leveraging SDEdit [34] as depicted in Figure 19, demonstrating the promise of achieving controllable 3D generation.
We provide more generated samples of unconditional and class-conditioned generation in Figure 20Figure 21 and Figure 22. The additional results demonstrate the strong capability of our model to create high-quality 3D assets with complex geometry and intricate textures.
Furthermore, we also provide an additional video in supplementary material, which intuitively illustrates our approach and visualizes the generated results.
A.4 Limitations
While GaussianCube represents a substantial step forward in developing an ideal representation for 3D content generation, it still has some limitations. Specifically, although the GaussianCube construction procedure is considerably more rapid than that of NeRF-based methods and can be executed in parallel, it still requires approximately 5 minutes to construct each object. This presents a challenge for scaling up training on extensive 3D datasets. In future work, we plan to investigate more time-efficient methods for GaussianCube construction. Additionally, akin to prior 2D diffusion models, our text-to-3D diffusion model encounters difficulties in presenting the specified number of objects within prompts as shown in Figure 13. To address this, we will look into enhancing the precision and controllability of 3D generation in the future.
A.5 Broader Impacts
The proposed GaussianCube enables high-quality 3D asset fitting with few parameters, which significantly simplifies the challenges of 3D generative modeling. Our diffusion model is capable of generating high-quality 3D assets of complex geometry and intricate textures while also accommodating a variety of conditional signals to steer the creating procedure. The strong capability of GaussianCube suggests its potential to serve as a versatile 3D representation for a variety of applications in future 3D research endeavors.
Like all generative models, particular caution is required when dealing with sensitive tasks involving human representations. Our avatar creation model is trained exclusively on a synthetic dataset [57] composed of large-scale 3D digital avatars which are generated through a graphics pipeline. We conceptualize digital avatars as analogous to those created by specialized 3D artists, rather than photorealistic human images. This strategy in selecting training data mitigates privacy and copyright issues that might arise from utilizing real human photo collections. Nevertheless, it is crucial to acknowledge that avatars generated by our model from real-world imagery could still be misused for spreading disinformation. As such, we advocate implementing rigorous safeguards and promoting responsible use of our technology other related ones to mitigate such risks.